# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@
thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work
@@
Ran Finkelstein    Richard Bing-Shiun Tsai    Xiangkai Sun California Institute of Technology, Pasadena, CA 91125, USA    Pascal Scholl California Institute of Technology, Pasadena, CA 91125, USA
@@
Enhancing the precision of measurements by harnessing entanglement is a long-sought goal in the field of quantum metrology 1, 2. Yet attaining the best sensitivity allowed by quantum theory in the presence of noise is an outstanding challenge, requiring optimal probe-state generation and readout strategies 3, 4, 5, 6, 7. Neutral atom optical clocks 8, leading systems for measuring time, have shown recent progress in terms of entanglement generation 9, 10, 11, but currently lack the control capabilities to realize such schemes. Here we show universal quantum operations and ancilla-based readout for ultranarrow optical transitions of neutral atoms. Our demonstration in a tweezer clock platform 12, 13, 14, 15, 9, 16 enables a circuit-based approach to quantum metrology with neutral atom optical clocks. To this end, we demonstrate two-qubit entangling gates with $>$99.35% fidelity via Rydberg interactions 17, 18, 15 and dynamical connectivity 19 for optical clock qubits, which we combine with local addressing 16 to implement universally programmable quantum circuits. Using this approach, we generate a near-optimal entangled probe state 1, 4, a cascade of Greenberger-Horne-Zeilinger (GHZ) states of different sizes, and perform dual-quadrature 5 GHZ readout. We also show repeated fast phase detection with non-destructive conditional reset of clock qubits and minimal dead time between repetitions by implementing ancilla-based quantum logic spectroscopy 20 (QLS) for neutral atoms. Finally, we extend this to multi-qubit parity checks and measurement-based Bell state preparation 21, 22, 23, 24. Our work lays the foundation for hybrid processor-clock devices with neutral atoms and more generally points to a future of practical applications for quantum processors linked with quantum sensors25.
@@
Figure 1: Universal quantum operations for tweezer clocks. a, We demonstrate a universal quantum processor based on optical clock qubits. Such a processor is designed to perform any computational task and at the same time is a highly sensitive clock. To realize such a device, we demonstrate high-fidelity entangling gates between optical clock qubits and circuits with dynamical reconfiguration. We combine this with local single-qubit rotations via single-site shifts, and local mid-circuit readout. b, In our experiment, we use tweezer-trapped ${}^{88}$r atoms, where qubits are based on the ultranarrow optical clock transition ${}^{1}\text{S}_{0}\leftrightarrow^{3}\text{P}_{0}$ and entangling operations are realized through transient excitation to a Rydberg state $n^{3}\text{S}_{1}$. We further utilize coherent superpositions of quantized motional states in tweezers to shelve protected clock qubits for mid-circuit detection of ancilla qubits. c, We identify two instances where these universal quantum operations can be used in optimizing quantum metrology: entangled state preparation for near-optimal probes, and entanglement-assisted readout with ancilla qubits, realizing quantum logic spectroscopy for neutral atoms.
@@
Control over quantum systems is evolving rapidly on many fronts. On the one hand, quantum processors – which use quantum mechanics to perform calculations – have matured into intermediate-scale systems with high-fidelity entangling operations and universal control capabilities 26, 27, 28, 29, 30, 19, 31. On the other hand, quantum sensors 25 and in particular optical atomic clocks 8 – which use quantum mechanics to measure time – have become both more accurate and precise, revealing new phenomena 32 and measuring relativistic effects at a laboratory scale 33, 34. It is thus a natural question, and a long-sought goal, how to utilize quantum information processing tools and entanglement to improve quantum sensors.
@@
Two main strategies are typically considered in this context 1. First, entanglement introduces correlations between probes, resulting in jointly enhanced sensitivity; unfortunately, this also entails jointly enhanced vulnerability to errors and reduced dynamic range, often nullifying or strongly reducing any effective gains 3, 35, 4, 36, 37. Schemes put forward to circumvent this naïve drawback use specific, finely-tuned entangled probe-state preparation and readout schemes 1, 4, 6. While recently one such scheme has been realized with trapped ions 7, there is so far no demonstration in neutral atom optical clocks, despite recent progress in entanglement generation 9, 10, 11.
@@
Second, entanglement with an ancillary system can be harnessed in non-destructive readout of a sensor to benefit metrology, e.g., logic spectroscopy of inaccessible probes as demonstrated with ions 20, sequential multi-round phase estimation 1, fast mid-circuit detection for multi-ensemble metrology 5, and shallow-circuit preparation of long-range entangled states 21, 22, 23, 24. Only very recently, ancilla-based readout for hyperfine qubits in neutral atoms has been shown 24, 38, but a demonstration for ultranarrow optical transitions – as needed for optical atomic clocks – is outstanding. Further, it is critical that ancilla-based readout can be performed repeatedly to fully realize its potential for metrology (and for quantum error correction 39), which is an outstanding challenge for neutral atoms in general.
@@
To demonstrate such complex entangled probe-state generation and repeated ancilla-based readout, we realize a scalable universal quantum processor with neutral atom qubits encoded in an ultranarrow optical transition (Fig. 1a). Importantly, the same experimental system can be utilized as a neutral atom optical clock 8, which currently forms the most stable frequency reference 33, 34. Our system specifically is a tweezer optical clock 12, 13, 14, 15, 9, 16, composed of neutral strontium-88 trapped in a tweezer array 40, 41 (Fig. 1b,c), which we now equip with universal quantum computing and repeated ancilla-based readout capabilities.
@@
In particular, we demonstrate high-fidelity entangling gates 27, 17, 18, 42 for optical clock qubits 15, achieving a controlled-Z (CZ) fidelity exceeding $0.9935(10)$ and generate entangled optical Bell pairs with a record fidelity. We further show optical coherence is preserved during qubit transport in tweezers, previously realized with hyperfine qubits 19, 43, enabling dynamical circuit reconfiguration. Combined with single-site control 16 and mid-circuit readout 44, this allows us to take a circuit-based approach to realize strategies for entangled state preparation and readout designed to improve metrology (Fig. 1c).
As part of this, we demonstrate simultaneous generation of a cascade of different-sized entangled GHZ states towards a quantum metrology scheme 1, 4 that circumvents the trade-off between entanglement-enhanced sensitivity and loss of dynamic range. This is particularly relevant in the context of state-of-the-art neutral atom optical clocks for which a loss of dynamic range implies an effective loss of stability, and where there is no expected gain, e.g., from fixed-sized GHZ states. Similarly, naive utilization of squeezed states is expected to yield only marginal gains 36, 37, 7. We also demonstrate ancilla-based mid-evolution readout, effectively realizing quantum logic spectroscopy 20 for neutral atoms, with numerous possible applications. First, this enables readout and reset with very low dead time, which could minimize the Dick effect, a current limitation of neutral atom optical clocks 8. Second, fully realizing the aforementioned quantum metrology scheme 1, 4 also requires implementing classically optimal clock interrogation 5, 4. While there is recent progress towards this end using dynamical decoupling 16, an advantageous avenue in terms of robustness to frequency noise consists of fast mid-evolution readout and reset, as demonstrated here. Third, we demonstrate ancilla-based readout for multi-qubit observables, which we use for a measurement-based generation of GHZ states, enabling in principle generation of large-scale GHZ states with shallow circuits 21, 22, 23, 24. These results more generally point to practical applications of quantum computers integrated in or connected with quantum sensors, as discussed in more detail in the outlook.
@@
High-fidelity entangling gates for optical qubits We begin by demonstrating high-fidelity entangling gates for the ultranarrow optical clock transition. The optical transition defines a qubit composed of the ground state of strontium-88 $\mathrm{{}^{1}S_{0}}$, denoted as $|0\rangle$, and the metastable state $\mathrm{{}^{3}P_{0}}$, denoted as $|1\rangle$ (Fig. 1b). To realize high-fidelity entangling operation, we apply a single, phase-modulated, Rydberg pulse for a CZ gate, which was shown to be time-optimal 42 and recently implemented in hyperfine and nuclear atomic qubits 17, 18. To simultaneously realize high-fidelity single-qubit rotations, we operate at a magnetic field of 450 G which balances between sufficiently high Rabi frequency on the clock transition (scaling linearly with the magnetic field 45) and sufficiently strong interactions (as the Rydberg interaction strength is magnetic field dependent, see Methods, Ext. Data Fig. 1). We prepare atoms in the $|1\rangle$ state and perform erasure-cooling 44 which prepares atoms close to the motional ground state.
Figure 2: High-fidelity entangling gates for optical clock qubits. a, Global randomized benchmarking (RB) circuit with a fixed number $\mathrm{N_{rep}}$ of random single-qubit gates $R_{C}$ and a variable number $\mathrm{N_{CZ}}$ of CZ gates. A final rotation $R_{f}$ is precomputed to maximize the return probability $P_{|11\rangle}$ in the absence of errors. b, By fitting the return probability as a function of $\mathrm{N_{CZ}}$, we infer an RB fidelity of $0.9935(10)$. We find this to be consistent with a full error model which predicts an RB fidelity of $0.9943(12)$ (shaded area denotes model uncertainty), accounting for both CZ errors and single-qubit errors. Without single-qubit errors, our model predicts a CZ fidelity of $0.9963(2)$, indicating the true experimental CZ fidelity might be underestimated by $2(1)\cdot 10^{-3}$ due to residual sensitivity of the RB sequence to single-qubit errors. c, A circuit for the creation and the characterization of the optical Bell state $|\Phi^{+}\rangle$. d, Parity oscillation with a measured bare parity contrast of $0.963^{+7}_{-10}$. Together with the population overlap (Methods), we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected). e, Array configuration depicted by average atomic fluorescence image. The interatomic separation is chosen to induce strong Rydberg interactions within a pair and avoid unwanted interactions between pairs.
@@
Figure 3: Simultaneous generation of a cascade of GHZ states and GHZ dual-quadrature readout. a, A quantum circuit for the simultaneous preparation and characterization of GHZ states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$, here $j=0,1,2$. b, Array reconfiguration steps depicted by average atomic fluorescence images at each array configuration (1), (2), (3), as marked in the circuit in a. Dashed arrows represent the atom movements between two array configurations. c, Simultaneous parity oscillations for $j=0,1,2$, from top to bottom, showing a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity. d, Simultaneous dual-quadrature parity readout of two copies of a 4-atom GHZ state realized through a $\pi/8$ phase shift on each atom in one copy, resulting in a collective $\pi/2$ phase shift before the final readout pulse.
We benchmark the performance of such gates by performing global two-qubit randomized benchmarking (RB) with single-qubit rotations 18, 17 drawn from the Pauli group. By design, the RB circuit in Fig. 2a results in maximal return probability of both atoms to the clock state ($P_{|11\rangle}$) in the case of perfect entangling gates and single-qubit rotations. Varying the number of CZ gates, $\mathrm{N_{CZ}}$, we infer an RB fidelity of $0.9935(10)$ by fitting the decay of the return probability as a function of $\mathrm{N_{CZ}}$ to an exponential trend (Fig. 2b).
The two-qubit RB circuit used here includes a fixed number of single-qubit pulses, to be minimally sensitive to single-qubit errors. However, we find that it has residual sensitivity to noise acting at the single-qubit level and thus likely underestimates the true CZ fidelity. Specifically, a detailed numerical model predicts an RB fidelity of $0.9963(2)$ including only two-qubit errors, which drops to $0.9943(12)$ if single-qubit errors are added. The model prediction, with both types of errors included (dashed line and shaded area in Fig. 2b), is consistent with the experimental RB result. Moreover, our model is independently found to be consistent with experimental results at the single-qubit level for a variety of test circuits (Methods), including single-qubit RB sequence with experimental fidelity of $0.9978(4)$ per $\pi/2$ pulse. The combination of experiment and consistent modeling results suggests that the true CZ fidelity is $2(1)\cdot 10^{-3}$ higher than $0.9935(10)$.
We note that independent of this correction, our result is only the second reported demonstration for neutral atoms generally, and the first in alkaline-earth atoms, to achieve two-qubit gate fidelity above 0.99 with Rydberg-based gates for long-lived qubits 17.
@@
As an important application of the CZ gate, we generate an array of Bell pairs with optical qubits (Fig. 2c). After initializing atoms in superposition states and applying a CZ-gate, a $\pi/4$ pulse rotates the resulting quantum state into the even-parity, metrologically relevant, Bell state $|\Phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$. For the state obtained in the experiment, we measure the population overlap with the ideal Bell state as well as drive parity oscillations (Fig. 2d) to infer the off-diagonal (coherence) terms 46. Overall, we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected). This is consistent with a CZ fidelity of 0.996 and single-qubit fidelity of 0.998 per $\pi/2$ pulse from our model prediction.
@@
Having demonstrated high-fidelity gates on optical clock qubits, we now turn to preparing larger entangled states and executing algorithms requiring more complex circuits with single-site addressing and dynamical reconfiguration. Specifically, we first demonstrate the preparation of a complex quantum register with metrological significance based on a cascade of entangled GHZ states. Such GHZ states enable, in principle, the best sensitivity allowed by quantum theory - the Heisenberg limit - with precision scaling as $1/N$, where $N$ is the atom number. However, for a single $N$-qubit GHZ state, while the sensitivity to phase variations is increased by a factor of $N$, the dynamic range is likewise reduced by a factor of $N$ 3, 4, 35. A similar trade-off exists for spin-squeezed entangled states 36, 37.
@@
Specifically, in current neutral atom optical atomic clocks, a smaller dynamic range also implies effectively reduced clock stability. Dynamic range here refers to the maximum interval over which phase excursions of a clock laser (the local oscillator) can be measured and corrected for. Since phase excursions grow with time, a small dynamic range means a shorter allowable interrogation time. As the precision of phase estimation is reduced with shorter interrogation time, a smaller dynamic range results in a loss of precision. Therefore, in practice, the nominally improved sensitivity of GHZ states is counteracted by the loss of dynamic range, and no improvement of clock stability from fixed-$N$ GHZ states is expected over optimal interrogation with uncorrelated atoms 3, 1, 4, 35. For spin-squeezed states, similarly, only moderate improvements are predicted 36, 37, 7.
@@
A proposed scheme 1, 4 to circumvent this trade-off between sensitivity and dynamic range requires the simultaneous preparation of a cascade of $M$ different groups of GHZ-like states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$ with $j=0,1...M-1$. Such a probe state enables Heisenberg-limited sensitivity, up to logarithmic corrections, while preserving the dynamic range. It thus enables a viable path to quantum-enhanced metrology under realistic conditions where local oscillator noise is the main limitation, as in current neutral atom optical clocks 47, 35. We note that these proposals are quantum-enhanced extensions of the classical optimal interrogation schemes 5, 48 that were recently experimentally shown 16, 49.
We realize a quantum circuit (Fig. 3a) for the required simultaneous preparation of GHZ states with $j=0,1,2$, utilizing our universal quantum processor aided by dynamical array reconfiguration (Fig. 3b). We note that this specific circuit is constrained in a one-dimensional system to produce GHZ states of up to four atoms. However, with further local operations that are available in our processor, one can in principle prepare arbitrary-size GHZ states with deeper circuits (Ext. Data Fig. 5). We also discuss a straightforward extension to two-dimensional arrays in the outlook, which would enable creating several copies of each GHZ state. We verify the simultaneous preparation of increasingly phase-sensitive states by varying the phase of a final $\pi/2$ pulse and measuring the parity signal $\mathcal{P}=\prod_{i}{Z_{i}}$ for each GHZ state. In Fig. 3c, we plot this signal against phase and find a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity, consistent with the 1:2:4 ratio expected.
@@
In this demonstration, only one quadrature of the optical phase signal is measured. Importantly, the dynamic range can be doubled if both phase quadratures are probed simultaneously, so-called dual-quadrature readout 4, 5, which has been realized recently with uncorrelated optical clock qubits 16, 49. With extended dynamic range, the maximal interrogation time is increased, and hence the sensitivity grows. We demonstrate such dual-quadrature readout for entangled states (Fig. 3d) by preparing two 4-atom GHZ states, and then imparting a phase rotation to only one through local phase shifts. This enables readout in the $y$-basis in parallel to readout in the $x$-basis. Here, a $\pi/8$ phase shift per atom ($\pi/2$ phase on the entangled four-atom register) is realized by moving 16 all four atoms by $\lambda/16\approx 44$ nm. We note such phase shifts could alternatively be realized by moving a single atom in the GHZ state by $\lambda/4$.
@@
An important question concerns the limitations on the maximum attainable GHZ state size in terms of achievable fidelity. We hypothesize that the phase noise of our clock laser is the dominant error source, which we test with experimental and numerical investigations. In the simultaneous GHZ state preparation scheme, we achieve a bare parity contrast of $0.52(3)$ for $N$=4 (Fig. 3c). To reduce the impact of clock laser phase noise, we prepare $N$=4 and $N$=2 GHZ states using a circuit with shorter idle time, i.e. wait time during which operations are done on other registers (by removing the single site operation $Z^{\frac{1}{4}}$ shown in Fig. 3a). With this sequence, we find an improved contrast of $0.68(3)$ and fidelity of $0.71(2)$ (Ext. Data Fig. 9). We note that the contrasts from both sequences are at or above the threshold of $0.51/\sqrt{N}=$ 0.5 required for a metrological gain 50, 2 at a fixed, short interrogation time. However, they do not exceed the threshold for the cascaded-GHZ scheme, which we find is higher (Methods).
@@
Motivated by this, we study fidelity as a function of the circuit idle time. We measure a decay in fidelity and a modification of the bitstring distribution (Ext. Data Fig. 9) which follows the prediction of our error model, consistent with our hypothesis that clock laser phase noise is the dominant error source. We further run a model using a frequency noise spectrum of a 26-mHz clock laser system 51, for which we predict a $N$=4 GHZ state fidelity $\approx$ 0.97 (Methods, Ext. Data Fig. 6). Thus, we expect significantly improved results with an upgraded clock laser system, reaching beyond the metrological gain threshold for the cascaded-GHZ scheme, even for larger $N$ (Methods, Ext. Data Fig. 7).

### v2

@@
thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work
@@
Ran Finkelstein    Richard Bing-Shiun Tsai    Xiangkai Sun    Pascal Scholl California Institute of Technology, Pasadena, CA 91125, USA
@@
Enhancing the precision of measurements by harnessing entanglement is a long-sought goal in the field of quantum metrology 1, 2. Yet attaining the best sensitivity allowed by quantum theory in the presence of noise is an outstanding challenge, requiring optimal probe-state generation and readout strategies 3, 4, 5, 6, 7. Neutral atom optical clocks 8, leading systems for measuring time, have shown recent progress in terms of entanglement generation 9, 10, 11, but currently lack the control capabilities to realize such schemes. Here we show universal quantum operations and ancilla-based readout for ultranarrow optical transitions of neutral atoms. Our demonstration in a tweezer clock platform 12, 13, 14, 15, 9, 16 enables a circuit-based approach to quantum metrology with neutral atom optical clocks. To this end, we demonstrate two-qubit entangling gates with $99.62\%$ fidelity via Rydberg interactions 17, 18, 15 and dynamical connectivity 19 for optical clock qubits, which we combine with local addressing 16 to implement universally programmable quantum circuits. Using this approach, we generate a near-optimal entangled probe state 1, 4, a cascade of Greenberger-Horne-Zeilinger (GHZ) states of different sizes, and perform dual-quadrature 5 GHZ readout. We also show repeated fast phase detection with non-destructive conditional reset of clock qubits and minimal dead time between repetitions by implementing ancilla-based quantum logic spectroscopy 20 (QLS) for neutral atoms. Finally, we extend this to multi-qubit parity checks and measurement-based, heralded, Bell state preparation 21, 22, 23, 24. Our work lays the foundation for hybrid processor-clock devices with neutral atoms and more generally points to a future of practical applications for quantum processors linked with quantum sensors25.
@@
Figure 1: Universal quantum operations for tweezer clocks. a, We demonstrate a universal quantum processor based on optical clock qubits. Such a processor is designed to perform any computational task and at the same time is a highly sensitive clock. To realize such a device, we demonstrate high-fidelity entangling gates between optical clock qubits and circuits with dynamical reconfiguration. We combine these with local single-qubit rotations via sub-wavelength atomic shifts and local mid-circuit readout. b, In our experiment, we use tweezer-trapped 88Sr atoms, where qubits are based on the ultranarrow optical clock transition ${}^{1}\text{S}_{0}\leftrightarrow^{3}\text{P}_{0}$, and entangling operations are realized through transient excitation to a Rydberg state $n^{3}\text{S}_{1}$. We further utilize coherent superpositions of quantized motional states in tweezers to shelve protected clock qubits for mid-circuit detection of ancilla qubits. c, We identify two instances where these universal quantum operations can be used in optimizing quantum metrology: entangled state preparation for near-optimal probes and entanglement-assisted readout with ancilla qubits, realizing quantum logic spectroscopy for neutral atoms.
@@
Control over quantum systems is evolving rapidly on many fronts. On the one hand, quantum processors – which use quantum mechanics to perform calculations – have matured into intermediate-scale systems with high-fidelity entangling operations and universal control capabilities 26, 27, 28, 29, 30, 19, 31, 32, 33. On the other hand, quantum sensors 25 and in particular optical atomic clocks 8 – which use quantum mechanics to measure time – have become both more accurate and precise, revealing new phenomena 34 and measuring relativistic effects at a laboratory scale 35, 36. It is thus a natural question, and a long-sought goal, how to utilize quantum information processing tools and entanglement to improve quantum sensors.
@@
Two main strategies are typically considered in this context 1: entangled state preparation and entanglement-assisted measurements. For the first, entanglement introduces correlations between probes, resulting in jointly enhanced sensitivity; unfortunately, this also entails jointly enhanced vulnerability to errors and reduced dynamic range, often nullifying or strongly reducing any effective gains 3, 37, 4, 38, 39. Schemes put forward to circumvent this apparent drawback use specific, finely-tuned entangled probe-state preparation and readout schemes 1, 4, 6. While recently one such scheme has been realized with trapped ions 7, there is so far no demonstration in neutral atom optical clocks, despite recent progress in entanglement generation 9, 10, 11.
@@
Second, entanglement with an ancillary system can be harnessed in non-destructive readout of a sensor to benefit metrology, e.g., logic spectroscopy of inaccessible probes as demonstrated with ions 20, sequential multi-round phase estimation 1, fast mid-circuit detection for multi-ensemble metrology 5, and shallow-circuit preparation of long-range entangled states 21, 22, 23, 24. Only very recently, ancilla-based readout for hyperfine qubits in neutral atoms has been shown 24, 40, but a demonstration for ultranarrow optical transitions – as needed for optical atomic clocks – is outstanding. Further, it is critical that ancilla-based readout can be performed repeatedly to fully realize its potential for metrology (and for quantum error correction 41), which is an outstanding challenge for neutral atoms in general.
@@
To demonstrate such complex entangled probe-state generation and repeated ancilla-based readout, we realize a scalable universal quantum processor with neutral atom qubits encoded in an ultranarrow optical transition (Fig. 1a). Importantly, the same experimental system can be utilized as a neutral atom optical clock 8, which currently forms the most stable frequency reference 35, 36. Our system specifically is a tweezer optical clock 12, 13, 14, 15, 9, 16, composed of neutral strontium-88 trapped in a tweezer array 42, 43 (Fig. 1b,c), which we now equip with universal quantum computing and repeated ancilla-based readout capabilities.
@@
In particular, we demonstrate high-fidelity entangling gates 27, 17, 18 for optical clock qubits 15, achieving a state-of-the-art controlled-Z (CZ) fidelity for neutral atoms of $0.9962(4)$ (averaged over two-qubit symmetric stabilzier states), and generate an array of optical-clock-transition Bell pairs with a record fidelity. We further show that optical coherence is preserved during qubit transport in tweezers, previously realized with hyperfine qubits 19, 44, enabling dynamical circuit reconfiguration. Combined with single-site control 16 and mid-circuit readout 45, this allows us to take a circuit-based approach to realize strategies for entangled state preparation and readout designed to improve metrology (Fig. 1c).
@@
High-fidelity entangling gates for optical qubits We begin by demonstrating high-fidelity entangling gates for the ultranarrow optical clock transition. The optical transition defines a qubit composed of the ground state of strontium-88 $\mathrm{{}^{1}S_{0}}$, denoted as $|0\rangle$, and the metastable state $\mathrm{{}^{3}P_{0}}$, denoted as $|1\rangle$ (Fig. 1b). To realize high-fidelity entangling operation, we apply a single phase-modulated Rydberg pulse for a CZ gate, which was shown to be time-optimal 46, 47 and recently implemented in hyperfine and nuclear atomic qubits 17, 18. To simultaneously realize high-fidelity single-qubit rotations, we operate at a magnetic field of 450 G which balances between sufficiently high Rabi frequency on the clock transition (scaling linearly with the magnetic field 48) and sufficiently strong interactions (as the Rydberg interaction strength is magnetic field dependent, see Methods, Ext. Data Fig. 1). We prepare atoms in the $|1\rangle$ state and perform erasure-cooling 45 which prepares atoms close to the motional ground state.
[…]
@@
Figure 3: Simultaneous generation of a cascade of GHZ states and GHZ dual-quadrature readout. a, A quantum circuit for the simultaneous preparation and characterization of GHZ states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$, here $j=0,1,2$. b, Array reconfiguration steps depicted by average atomic fluorescence images at each one-dimensional array configuration (1), (2), (3), as marked in the circuit in a. Dashed arrows represent atom movements performed between two sequential array configurations. c, Simultaneous parity oscillations for $j=0,1,2$, from top to bottom, showing a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity. d, Simultaneous dual-quadrature parity readout of two copies of a 4-atom GHZ state realized through a $\pi/8$ phase shift on each atom in one copy, resulting in a collective $\pi/2$ phase shift before the final readout pulse.
We characterize the performance of such gates via a randomized circuit (Fig. 2a) designed to benchmark the CZ gate fidelity averaged over the symmetric stabilizer states, which are the set of states accessible with global single-qubit Clifford gates and two-qubit CZ gates (Methods, Ext. Data Fig. 3). Varying the number of CZ gates $\mathrm{N_{CZ}}$, we infer a CZ fidelity of $0.9962(4)$ by fitting the decay of the return probability as a function of $\mathrm{N_{CZ}}$ to an exponential trend (Fig. 2b). The quoted fidelity is corrected for false contribution from leakage out of the qubit manifold during the gate (Methods). We numerically find this gate benchmark to be within $5\%$ of the infidelity averaged over Haar random states for our gate parameters. Our result sets a new state-of-the-art fidelity in Rydberg-based two-qubit entangling gates for long-lived qubits 17, 49.
@@
As an important application of the CZ gate, we generate an array of Bell pairs with optical clock qubits (Fig. 2c). After initializing atoms in superposition states and applying a CZ-gate, a $\pi/4$ pulse rotates the state into the even-parity, metrologically relevant, Bell state $|\Phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$. For the state obtained in the experiment, we measure the population overlap with the ideal Bell state as well as drive parity oscillations (Fig. 2d) to infer the off-diagonal (coherence) terms 50. Overall, we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected) which is limited, beyond the CZ fidelity, by the finite fidelity of clock $\pi/2$ pulses and decoherence during the circuit (Methods).
@@
Having demonstrated high-fidelity gates on optical clock qubits, we now turn to preparing larger entangled states and executing algorithms requiring more complex circuits. Specifically, we first demonstrate the preparation of a complex quantum register with metrological significance based on a cascade of entangled GHZ states. Such GHZ states enable, in principle, the best sensitivity for linear observables allowed by quantum theory - the Heisenberg limit - with precision scaling as $1/N$, where $N$ is the atom number. However, for a single $N$-qubit GHZ state, while the sensitivity to phase variations is increased by a factor of $N$, the dynamic range is likewise reduced by a factor of $N$ 3, 4, 37. A similar trade-off exists for spin-squeezed entangled states 38, 39.
@@
Specifically, in current neutral atom optical atomic clocks, a smaller dynamic range also implies effectively reduced clock stability. Dynamic range here refers to the maximum interval over which phase excursions of a clock laser (the local oscillator) can be unambiguously measured and corrected for. Since phase excursions grow on average with time, a smaller dynamic range means a shorter allowable interrogation time. As the precision of phase estimation is reduced with shorter interrogation time, a smaller dynamic range results in a loss of precision. Therefore, in practice, the nominally improved sensitivity of GHZ states is counteracted by the loss of dynamic range, and no improvement of clock stability from fixed-$N$ GHZ states is expected over optimal interrogation with uncorrelated atoms 3, 1, 4, 37. For spin-squeezed states, similarly, only moderate improvements are predicted 38, 39, 7.
@@
A proposed scheme 1, 4 to circumvent this trade-off between sensitivity and dynamic range requires the simultaneous preparation of a cascade of $M$ different groups of GHZ-like states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$ with $j=0,1...M-1$. Such a probe state enables Heisenberg-limited sensitivity, up to logarithmic corrections, while preserving the dynamic range. It thus enables a viable path to quantum-enhanced metrology under realistic conditions where local oscillator noise is the main limitation, as in current neutral atom optical clocks 51, 37. We note that these proposals are quantum-enhanced extensions of the classical optimal interrogation schemes 5, 52 that were recently experimentally shown 16, 53.
We consider different circuit realizations for generating the required simultaneous preparation of GHZ states with $j=0,1,2$, utilizing our universal quantum processor aided by dynamical array reconfiguration. While large GHZ states can be prepared with circuits applying extensive local control (Ext. Data Fig. 5), here we realize a circuit compiled with global control up to its final layer (Fig. 3a). This choice is guided by minimizing idle time added by local operations, during which laser-atom dephasing occurs. The local control for single-qubit rotations is achieved through local phase shifts imparted via sub-wavelength shifts in atomic position 16, and the connectivity for two-qubit gates is realized by mid-circuit array reconfiguration (Fig. 3b). We verify the simultaneous preparation of increasingly phase-sensitive states by varying the phase of a final $\pi/2$ pulse and measuring the parity signal $\mathcal{P}=\prod_{i}{Z_{i}}$ for each GHZ state. In Fig. 3c, we plot this signal against phase and find a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity, consistent with the 1:2:4 ratio expected.
@@
This demonstration measures only one quadrature of the optical phase signal. Importantly, the dynamic range can be doubled if both phase quadratures are probed simultaneously, so-called dual-quadrature readout 4, 5, 16, 53. We demonstrate such dual-quadrature readout for entangled states (Fig. 3d) by preparing two 4-atom GHZ states and then imparting a phase rotation to only one through local phase shifts. This enables readout in the ${\hat{Y}}$-basis in parallel to readout in the ${\hat{X}}$-basis. Here, a $\pi/8$ phase shift per atom ($\pi/2$ phase on the entangled 4-atom register) is realized by moving 16 all four atoms by $\lambda/16\approx 44$ nm. We note that such phase shifts could alternatively be realized by moving a single atom in the GHZ state by $\lambda/4$.
@@
An important question concerns the limitations on the maximum attainable GHZ state size in terms of achievable fidelity. We hypothesize that the phase noise of our clock laser is the dominant error source, which we test with experimental and numerical investigations. In the simultaneous GHZ state preparation scheme, we achieve a bare parity contrast of $0.52(3)$ for $N$=4 (Fig. 3c). To reduce the impact of clock laser phase noise, we prepare $N$=4 and $N$=2 GHZ states using a circuit with shorter idle time, i.e. wait time during which operations are done on other registers (by removing the single site operation $Z^{\frac{1}{4}}$ shown in Fig. 3a). With this sequence, we find an improved contrast of $0.68(3)$ and fidelity of $0.71(2)$ (Ext. Data Fig. 9). We note that the contrasts from both sequences are at or above the threshold of $0.51/\sqrt{N}=$ 0.5 required for a metrological gain 54, 2 at a fixed, short interrogation time. However, they do not exceed the higher threshold for the cascaded-GHZ scheme (Methods).
@@
Motivated by this, we study fidelity as a function of the circuit idle time. We measure a decay in fidelity and a modification of the bitstring distribution (Ext. Data Fig. 9) which follows the prediction of our error model, consistent with our hypothesis that clock laser phase noise is the dominant error source. We further run a model using a frequency noise spectrum of a 26-mHz clock laser system 55, for which we predict a $N$=4 GHZ state fidelity $\approx$ 0.97 (Methods, Ext. Data Fig. 6). Thus, we expect significantly improved results with an upgraded clock laser system, surpassing the metrological gain threshold for the cascaded-GHZ scheme, even for larger $N$ (Methods, Ext. Data Fig. 7).